# Sparse Mixture of Experts as Unified Competitive Learning

### v1

@@ 4.1 Experiment Setting
Without Training.  We evaluate our approach on three state-of-the-art Sparse Mixture of Experts (MoE) models: (1) OLMoE-1B-7B Muennighoff et al. (2024), which has 7B parameters, 16 layers, and 64 experts per layer; (2) Qwen1.5-MoE-A2.7B Team (2024), comprising 7B parameters, 24 layers, and 60 experts per layer; and DeepSeekMoE-16B which consists of 16B parameters, 28 layers, and 64 experts per layer. We evaluate these models on the Massive Text Embedding Benchmark (MTEB) without additional fine-tuning. Additionally, we compare methods both with and without prompts, including PromptEOL Jiang et al. (2024c). Inspired by  Li and Zhou (2024), we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  Li and Zhou (2024), which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
Pre-training.  To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE Jiang et al. (2024b), SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE Chi et al. (2022b), and StableMoE Dai et al. (2022a), as well as the Expert Choice approach Zhou et al. (2022a) for pre-training and fine-tuning tasks. We follow the approach of Chen et al. (2023b) and use a base Transformer-XL Dai et al. (2019a) with four decoder layers. We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in Chen et al. (2023b).
Fine-tuning.  We fine-tune the pre-trained weights for text classification tasks, including SST-2 Socher et al. (2013), SST-5 Socher et al. (2013), IMDB Maas et al. (2011), and BANKING77 Casanueva et al. (2020). More implementation details and additional results are provided in the Appendix A.
4.2 Without Training
@@
Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL Jiang et al. (2023) (denoted as "with prompts" for brevity), as shown in Table 1, and (2) without prompts, as shown in Table 2. Detailed results for datasets under each task type are provided in Appendix A. USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
Figure 1 illustrates the effectiveness of our method across various MTEB tasks using the OLMoE-1B-7B model, in comparison to both the Expert Choice and Token Choice approaches. For tasks evaluated with prompts, USMoE achieves the highest average performance across models, showing notable improvements of 4.2%, 9.1%, and 6.6% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, without requiring any additional training, as detailed in Table 1. Notably, DeepSeekMoE-16B demonstrates a significant improvement from 50.1% (Token Choice) to 64.3% (USMoE), reflecting a 14.2% gain in the STS task.
For tasks evaluated without prompts, USMoE proves even more effective at enhancing the Token Choice approach, delivering notable gains of 9.3%, 7.0%, and 8.0% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, across MTEB tasks without additional training. Specifically, OLMoE-1B-7B achieves a remarkable improvement from 24.1% (Token Choice) to 47.5% (USMoE), representing a 23.4% gain. This trend persists across Qwen1.5-MoE and OLMoE, where USMoE consistently outperforms both the Token Choice and Expert Choice approaches.
@@
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 59.2
Clustering 16.2 24.8 26.9 23.2 30.5
PairClassification 53.5 62.0 58.9 66.0 66.8
Reranking 41.7 51.3 51.0 53.2 54.7
STS 49.4 63.5 44.2 67.8 71.1
Summarization 25.6 28.9 29.7 30.4 30.9
Average 38.3 48.0 44.5 48.7 52.2
Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.2
Clustering 14.3 34.2 29.2 30.1 35.8
PairClassification 51.9 60.5 56.0 60.3 63.4
Reranking 41.0 46.6 45.0 51.1 53.7
STS 48.3 50.1 50.0 64.3 66.3
Summarization 27.0 23.0 21.9 27.3 48.4
Average 38.6 45.4 39.6 47.9 54.5
DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 57.3
Clustering 17.8 29.0 20.3 28.5 31.9
PairClassification 57.4 59.8 53.8 63.3 65.3
Reranking 43.8 45.7 40.9 50.6 52.1
STS 52.8 49.0 37.1 63.4 66.0
Summarization 29.1 24.4 25.7 29.2 30.9
Average 41.6 44.0 38.9 48.0 50.6
@@
Table 1: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL Jiang et al. (2023). The best result for each row is highlighted in bold.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 47.0
Clustering 13.7 14.7 12.0 14.5 16.9
PairClassification 45.3 39.1 35.5 45.7 50.1
Reranking 37.5 37.4 35.3 39.5 43.1
STS 39.9 24.1 18.2 39.9 47.5
Summarization 28.4 20.9 21.1 29.8 30.8
Average 34.3 29.9 27.8 35.2 39.2
Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 51.2
Clustering 13.6 27.4 23.2 25.2 26.7
PairClassification 45.9 46.9 43.4 51.5 53.1
Reranking 39.6 45.3 41.6 48.5 48.2
STS 38.8 38.0 35.6 51.8 54.8
Summarization 28.3 13.4 15.1 31.2 29.7
Average 35.0 36.9 30.7 42.6 43.9
DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 46.8
Clustering 13.4 18.1 13.5 17.8 21.9
PairClassification 45.5 40.9 37.1 46.1 51.9
Reranking 38.5 38.9 35.1 42.2 45.7
STS 37.7 26.3 23.3 40.2 45.9
Summarization 24.9 22.0 18.5 24.4 28.4
Average 33.9 32.1 28.7 35.9 40.1
@@
Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
[…]
SMoE-DR 1.56 1.56 58.37 93.17
XMoE 1.21 1.28 30.34 58.33
StableMoE 1.20 1.28 29.97 58.25
[…]
4.3 Pre-training Result
Base models training. Table 3 summarizes the pre-training results across four datasets (enwik8, text8, WikiText-103, and One Billion Words). USMoE consistently outperforms the expert choice baseline and token choice routing methods such as XMoE Chi et al. (2022a) and StableMoE Dai et al. (2022b) on all four datasets. The strength of USMoE lies in its inference efficiency, achieved by leveraging fewer experts. Notably, on text8, USMoE surpasses SMoE while utilizing only one and a half experts. Additionally, it outperforms SMoE-Dropout (which employs two experts) on One Billion Words, lowering perplexity from 93.17 to 57.55 with the same reduced expert count. Furthermore, with only one and a half experts, USMoE reduces FLOPs by 14% compared to SMoE and SMoE-Dropout, which rely on two experts, all while maintaining strong performance.
Transformer-XL(420M) WikiText-103 lm1b
[…]
1 31.70 35.52 25.48 58.65 65.43 56.90
2 22.42 23.30 22.06 44.56 43.39 40.53
4 23.57 23.60 22.65 45.52 43.70 40.90
8 24.20 24.37 22.88 46.36 44.22 43.24
4.4 Fine-tuning Result
[…]
SMoE-DR 78.6 34.4 83.5 66.7
XMoE 76.7 35.3 83.3 67.4
StableMoE 77.7 34.3 83.9 60.8
[…]
Fine-tuning.  We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  5, using Transformer-XL pre-trained on enwik8. Overall, USMoE consistently achieves higher accuracy compared to other baselines across all datasets. The results demonstrate that our method is not only effective for pre-training tasks but also performs effectively on existing pre-trained models.
4.5 Ablation Studies
4.5.1 Competitive Learning Strategy Comparison
@@
Table 6: Competitive Learning Strategy comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.
The Unified Competitive Mechanism is implemented using two approaches: (1) a sequence-based method that compares all tokens within a sequence (referred to as "USMoE-Sequence") and (2) a batch-based method that compares all tokens within a batch or mini-batch (referred to as "USMoE-Batch"). We evaluate both approaches on the Classification task, with results presented in Table 6. The findings indicate that both methods outperform the Expert Choice and Token Choice approaches, demonstrating the effectiveness of our method. Notably, the sequence-based approach achieves superior performance in the Classification task, as it ensures that no important tokens are missed within a sequence - an assurance that the batch/mini-batch implementation may not always provide.
4.5.2 Robustness to the controlling factor

### v2

@@ A.4.2 Training-free additional results
In this section, we compare methods both with and without prompts, including PromptEOL [29]. Inspired by  [34], we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  [34], which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 61.4
Clustering 16.2 24.8 26.9 23.2 32.1
PairClassification 53.5 62.0 58.9 66.0 68.9
Reranking 41.7 51.3 51.0 53.2 55.1
STS 49.4 63.5 44.2 67.8 71.1
Summarization 25.6 28.9 29.7 30.4 30.5
Average 38.3 48.0 44.5 48.7 53.2
Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.7
Clustering 14.3 34.2 29.2 30.1 37.5
PairClassification 51.9 60.5 56.0 60.3 66.6
Reranking 41.0 46.6 45.0 51.1 56.8
STS 48.3 50.1 50.0 64.3 69.0
Summarization 27.0 23.0 21.9 27.3 31.0
Average 38.6 45.4 39.6 47.9 53.4
DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 60.4
Clustering 17.8 29.0 20.3 28.5 32.8
PairClassification 57.4 59.8 53.8 63.3 67.9
Reranking 43.8 45.7 40.9 50.6 52.4
STS 52.8 49.0 37.1 63.4 68.1
Summarization 29.1 24.4 25.7 29.2 30.7
Average 41.6 44.0 38.9 48.0 52.0
Table 5: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 51.3
Clustering 13.7 14.7 12.0 14.5 21.0
PairClassification 45.3 39.1 35.5 45.7 61.7
Reranking 37.5 37.4 35.3 39.5 46.5
STS 39.9 24.1 18.2 39.9 59.6
Summarization 28.4 20.9 21.1 29.8 31.8
Average 34.3 29.9 27.8 35.2 45.3
Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 52.2
Clustering 13.6 27.4 23.2 25.2 29.9
PairClassification 45.9 46.9 43.4 51.5 56.2
Reranking 39.6 45.3 41.6 48.5 53.1
STS 38.8 38.0 35.6 51.8 60.7
Summarization 28.3 13.4 15.1 31.2 40.0
Average 35.0 36.9 30.7 42.6 48.7
DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 49.4
Clustering 13.4 18.1 13.5 17.8 21.0
PairClassification 45.5 40.9 37.1 46.1 53.5
Reranking 38.5 38.9 35.1 42.2 45.7
STS 37.7 26.3 23.3 40.2 51.2
Summarization 24.9 22.0 18.5 24.4 29.9
Average 33.9 32.1 28.7 35.9 41.8
@@
In this section, inspired by  [34], we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B [37]:, DeepSeekMoE-16B  [12], Qwen1.5-MoE-A2.7B [52]. We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  [38], which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization. Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization. Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL [28] (denoted as "with prompts" for brevity), as shown in Table 5, and (2) without prompts, as shown in Table 6. Detailed results for datasets under each task type are provided in Appendix A. USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
@@
We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  7), Qwen1.5-MoE-A2.7B (Table 8), and DeepSeekMoE-16B (Table  9). Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification OLMoE Emotion None 24.1 24.5 26.3 25.1 35.8
Prompt 27.6 49.9 49.0 44.5 54.8
Toxic None 51.9 58.9 59.9 51.9 62.8
Prompt 52.3 65.2 61.3 53.4 67.6
Tweet None 47.7 46.8 48.3 48.4 55.2
Prompt 49.5 58.0 58.4 57.2 61.7
Clustering OLMoE Medrxiv None 15.0 17.6 14.8 17.4 20.6
Prompt 15.8 23.9 27.7 22.0 28.2
20Groups None 12.4 11.8 9.2 11.5 21.3
Prompt 16.7 25.7 26.2 24.4 36.0
Pair Classification OLMoE SemEval None 43.6 35.8 31.3 43.6 50.2
Prompt 45.7 46.7 40.9 53.8 55.4
URLCorpus None 47.0 42.4 39.7 47.8 73.2
Prompt 61.4 77.4 76.9 78.2 82.3
Reranking OLMoE Ask None 41.3 41.0 39.0 41.4 47.1
Prompt 43.4 51.9 49.9 50.2 51.6
SciDocs None 45.5 46.3 46.9 50.8 59.1
Prompt 53.6 69.6 73.1 75.1 77.2
StackOver None 25.8 24.8 20.1 26.4 33.2
Prompt 28.1 32.5 30.0 34.3 36.6
STS OLMoE Biosses None 39.3 13.6 7.7 29.7 64.0
Prompt 51.2 61.8 67.6 70.2 75.2
SickR None 50.3 46.3 26.4 53.0 58.9
Prompt 51.9 65.7 37.6 66.1 66.5
STS12 None 40.1 8.6 11.1 37.8 58.2
Prompt 51.3 53.8 37.5 63.6 66.4
STS13 None 40.5 21.1 18.2 43.4 61.5
Prompt 52.5 66.5 40.4 72.7 76.4
STS14 None 29.5 13.4 13.3 31.7 52.9
Prompt 41.1 56.8 33.9 64.2 68.2
STS15 None 30.8 27.8 22.5 33.3 63.2
Prompt 46.4 69.3 38.4 66.4 72.5
STS16 None 46.5 38.9 28.9 45.8 60.7
Prompt 52.4 70.1 49.4 68.3 71.8
STSBen None 42.2 23.4 17.5 44.5 57.4
Prompt 48.6 63.6 48.9 70.7 72.1
Summarization OLMoE Medrxiv None 28.4 20.9 21.1 29.8 31.8
Prompt 25.6 28.9 29.7 30.4 30.5
@@
Table 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB tasks using the OLMoE model. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 37.1
Prompt 37.0 48.5 47.4 47.2 51.3
Toxic None 53.0 61.1 21.3 52.9 61.4
Prompt 53.4 64.5 19.8 54.1 65.5
Tweet None 51.1 55.9 25.1 55.9 58.2
Prompt 56.1 61.1 38.6 60.7 62.3
Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 25.3
Prompt 14.2 24.6 19.8 21.8 27.9
20Groups None 12.0 31.5 25.1 27.4 34.4
Prompt 14.4 43.8 38.6 38.4 47.0
Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 47.5
Prompt 47.0 52.4 46.3 52.4 57.7
URLCorpus None 49.8 54.9 52.1 60.6 64.8
Prompt 56.7 68.7 65.8 68.2 75.5
Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 49.1
Prompt 43.3 48.3 49.1 49.5 52.8
SciDocs None 49.6 60.6 55.3 67.0 71.6
Prompt 50.9 60.1 55.8 68.7 73.0
StackOver None 26.2 29.5 26.2 31.1 38.5
Prompt 28.8 31.3 30.2 35.2 44.5
STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 66.6
Prompt 55.1 55.8 48.5 68.4 74.5
SickR None 51.0 55.5 40.4 61.0 63.5
Prompt 50.2 59.7 51.1 64.3 69.1
STS12 None 40.2 16.9 18.6 46.3 52.3
Prompt 49.3 25.0 31.8 59.2 62.5
STS13 None 38.1 42.9 44.2 56.7 67.3
Prompt 53.3 57.5 54.6 73.4 75.6
STS14 None 28.1 26.5 25.6 45.4 53.7
Prompt 40.4 38.8 40.7 60.0 64.8
STS15 None 34.8 40.5 38.4 46.1 56.1
Prompt 40.7 52.3 54.2 58.8 66.8
STS16 None 47.6 51.0 48.1 58.1 64.4
Prompt 51.6 64.2 65.1 65.7 68.7
STSBen None 37.0 37.7 34.7 50.9 61.6
Prompt 45.6 47.8 54.5 64.5 70.0
Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 40.0
Prompt 27.0 23.0 21.9 27.3 31.0
Figure 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 31.3
Prompt 37.9 48.3 46.1 46.4 52.6
Toxic None 53.3 60.4 58.1 53.1 61.7
Prompt 53.1 62.4 62.5 53.6 67.5
Tweet None 51.0 51.9 49.5 52.6 55.2
Prompt 54.9 58.4 57.5 58.9 61.0
Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.7
Prompt 17.0 25.7 20.9 24.0 27.9
20Groups None 11.7 13.2 9.7 13.7 16.2
Prompt 18.6 32.3 19.8 33.0 37.6
Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 44.6
Prompt 48.4 47.2 46.6 51.3 55.7
URLCorpus None 46.4 41.7 41.6 48.6 62.4
Prompt 66.5 72.4 61.1 75.4 80.0
Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 44.9
Prompt 43.5 43.8 44.7 46.9 50.6
SciDocs None 48.2 50.6 44.7 57.1 61.9
Prompt 58.3 65.6 55.3 72.6 72.9
StackOver None 25.7 24.9 20.4 27.3 30.2
Prompt 29.7 27.6 22.6 32.3 33.6
STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 56.6
Prompt 47.0 40.1 41.5 57.6 67.2
SickR None 50.4 47.4 29.4 53.1 60.1
Prompt 56.0 61.9 38.7 65.8 68.0
STS12 None 44.0 4.3 13.9 45.0 48.6
Prompt 57.8 31.0 28.4 64.0 65.9
STS13 None 36.0 28.4 27.5 41.1 50.7
Prompt 55.3 56.0 41.2 70.9 76.1
STS14 None 25.4 12.0 13.0 28.2 41.2
Prompt 44.9 41.0 31.1 58.6 65.1
STS15 None 34.8 33.9 25.6 38.7 45.0
Prompt 49.7 46.5 33.0 58.5 64.1
STS16 None 44.9 34.4 33.1 46.9 55.2
Prompt 56.7 58.0 44.2 64.5 67.4
STSBen None 36.6 18.3 15.8 42.1 51.9
Prompt 54.9 57.7 39.0 67.8 71.0
Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 29.9
Prompt 29.1 24.4 25.7 29.2 30.7
@@
Table 10: Comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.